OpenStudy (anonymous):

solve the equation for p S=p/3+7q

OpenStudy (anonymous):

retype

OpenStudy (anonymous):

retype what?

OpenStudy (anonymous):

cant understand ur ques

OpenStudy (anonymous):

S= p/3 + 7q solve for p

OpenStudy (anonymous):

\[S=\frac{p}{3}+7q\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[S-7q=\frac{p}{3}\] \[3(S-7q)=p\]

OpenStudy (anonymous):

so is it going to be p= -21q + 3S

OpenStudy (anonymous):

sure